OpenStudy (anonymous):

Question 13 (Multiple Choice Worth 5 points) [9.05] What are the exact solutions of x2 − x − 4 = 0? x = the quantity of negative 1 plus or minus the square root of 15 all over 2 x = the quantity of 1 plus or minus the square root of 15 all over 2 x = the quantity of 1 plus or minus the square root of 17 all over 2 x = the quantity of negative 1 plus or minus the square root of 17 all over 2

OpenStudy (rob1525):

lol ;)

OpenStudy (anonymous):

I think it be D but its not right to me can you help me

OpenStudy (rob1525):

so, lets use the difference of squares method (also know as the pathagorian therom) to solve this.

OpenStudy (rob1525):

\[(-b +or- \sqrt{b^2-4*ac})/2a\]

OpenStudy (anonymous):

whats next

OpenStudy (rob1525):

so the equation us have has the form a^2-b-c=0

OpenStudy (rob1525):

sooo....\[-1+or-(\sqrt{-1^2-4*1*-4})/2*1\]

OpenStudy (anonymous):

I don't know how to do this sorry

OpenStudy (rob1525):

solve for -1^2-4*1*-4 first...k

OpenStudy (anonymous):

I am lost sorry

OpenStudy (rob1525):

okay sorry... what does -1^2*4*1*-4=?

OpenStudy (rob1525):

opps I meant -1^2-4*1*-4=?

OpenStudy (anonymous):

how do i solve for that

OpenStudy (rob1525):

-1^2=1 right-4*1which is -4*-4 which =16...so ur left with 17 right?

OpenStudy (anonymous):

yea

OpenStudy (rob1525):

-1^2 means negative one to the second power. you know that ya?

OpenStudy (anonymous):

2 right

OpenStudy (rob1525):

-1x -1 = -1^2 = 1

OpenStudy (anonymous):

oh ok

OpenStudy (rob1525):

k

OpenStudy (anonymous):

so the answer is C right

OpenStudy (rob1525):

YOUUUUUU GOT IT... NICE JOB.

OpenStudy (anonymous):

cool one more thing can you help me on this Question 14 (Multiple Choice Worth 5 points) [9.05] What are the approximate solutions of 2x2 − 7x = 3 rounded to the nearest hundredth? No real solutions x ≈ −0.77 and x ≈ 7.77 x ≈ −0.39 and x ≈ 3.89 x ≈ −3.89 and x ≈ 0.39

OpenStudy (anonymous):

I think its A but not sure lol

OpenStudy (rob1525):

k..use the pathagorian therom to solve this one too..but u will need a calc for this one

OpenStudy (rob1525):

the 3 on the right side of the equation is confusing u hu?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

will it B

OpenStudy (rob1525):

so just subtract 3 from both sides so that the right side is equal to 0... like this \[2x^2-7x-3=3-3 -> 2x^2-7x-3=0\] see,

OpenStudy (rob1525):

now put it in your calc \[-1*-7 +or- (\sqrt{-7^2-4*2*-3})/2*2\]

OpenStudy (rob1525):